A single-lane street 10 ft wide goes through a semicircular tunnel with radius 9 ft. How high is the tunnel at the edge of each lane? Roun off to 2 decimal places.

1
Expert's answer
2021-12-09T16:00:45-0500

Solution;

"R=9ft"

"x=\frac{10}{2}=5ft"

But;

"x^2+y^2=r^2"

"y^2=9^2-5^2"

"y=\sqrt{56}=7.48ft"